OpenStudy (anonymous):

HELP PLEASE! Place the indicated product in the proper location on the grid. (1 - 7x)(1 + 9x)

OpenStudy (anonymous):

Heres an example, use this to help: Multiply each term in one bracket by each in the other: (I assume x2 means x^2 / x squared) (x^2 + 3x + 1)(x^2 + x + 2) = x^4 + x^3 + 2x^2 + 3x^3 + 3x^2 + 6x + x^2 + x +2 Now collect all terms: = x^4 + 4x^3 + 6x^2 + 7x + 2

OpenStudy (anonymous):

O_O